A spherical metal shell A of radius RA and a solid metal sphere B of radius RB (A) are kept far apart and each is given charge ‘+Q’, Now they are connected by a thin metal wire Then
 
(a) EAinside = 0
(b) QA > QB
(c) s A/ sB = RB/RA
(d) EAonsurface < EBonsurface

Grade:6

1 Answers

Navjyot Kalra
9 years ago

(a,b,c,d)

Electric field inside a spherical metallic shell with charge on the surface is always zero. Therefore option [a] is

correct.

When the sells are connected with a thin metal wire then electric potentials will be equal, say V

? 1/ 4p ?0 QA/RA = 1 / 4 p ?0 QB­/RB = V

As RA > RB therefore QA > A­B option [b] is also correct.

As sA/sB QA/4pR2A/QB/4pR2B = R2B / R2A x 4p ?0 RAV / 4p ?0 RAV

? sA/sB = RB / RA

Option (c) is also correct

Also EA = sA/ sB = RB/RA < 1 ? EA < EB

Option (d) is also correct